# Coupling Magnons to an Opto-Electronic Parametric Oscillator

### v1

@@ IV Analytical Modeling
$+\sum_{m}\Big(\xi\,e^{+2\pi if_{p}t}\,a_{+m}^{\dagger}a_{-m}^{\dagger}+\xi^{*}\,e^{-2\pi if_{p}t}\,a_{+m}a_{-m}\Big)\,$, (5)
@@
with $\omega_{a}\equiv 2\pi f_{p}/2$ and $\Delta_{a}\equiv\mathrm{FSR}/2$. Since each OEPO mode acts as an independent drive of magnetization dynamics, it is convenient to replace the pair-creation term in Eq. (5) with an equivalent classical drive on each mode and relabel $\omega_{\pm m}\to\omega_{\sigma m}$ for compactness, with $\sigma=\pm$ denoting the original sideband index. The Hamiltonian in Eq. (5) then becomes:
@@
$+i\xi\sum_{m,\sigma=\pm}\left(e^{-i\omega_{\sigma m}t}a_{\sigma m}^{\dagger}-\mathrm{h.c.}\right)\,.$ (7)
@@
with $g_{T}=\sqrt{g_{am}^{2}+g^{2}}$, $\kappa_{\!\mathrm{eff}}=(\kappa g_{am}^{2}+\gamma_{p}g^{2})/g_{T}^{2}$, $\kappa_{\!\mathrm{dark}}=(\kappa g^{2}+\gamma_{p}g_{am}^{2})/g_{T}^{2}$, and $\varepsilon=(\kappa-\gamma_{p})g\,g_{am})/g_{T}^{2}$. Equations (18)–(20) show that, while the dark mode $d_{\sigma m}$ couples dissipatively to $b_{\sigma m}$, it remains decoupled from the magnon mode $c_{0}$, and it is therefore invisible to the FMR probe. In the near-resonant, small-contrast limit $\lvert\kappa-\gamma_{p}\rvert\ll\kappa+\gamma_{p}$, the bright mode is well approximated as a single lossy channel with effective linewidth $\kappa_{\mathrm{eff}}$ that hybridizes with $c_{0}$ at the collective rate $g_{T}$; the corresponding two-mode cooperativity is then rewritten as $\mathcal{C}=\frac{4g_{T}^{2}}{\kappa_{\!\mathrm{eff}}\,\gamma_{m}}$. Fourier transforming Eqs. (14)–(16) and solving for the MS mode response, we find
@@ V Conclusion
Third, by exploiting a dual-nonlinearity scheme, the OEMPO achieves a magnon-photon mode hybridization through combining the optoelectronic nonlinearity at the mixer (parametric pump) and the magnonic nonlinearity at the YIG (three-magnon scattering), offering a unique testbed for exploring coupled nonlinear phenomena at the hybrid interfaces among electronic, optical, and magnetic platforms. In particular, magnon modes under confined geometries or subject to nonlinear mixing processes can also give rise to a magnonic frequency comb that is controllable by external magnetic field Hula et al. (2022); Wang et al. (2024b); Christy et al. (2025). This way it is possible to achieve mode hybridization between two interfering frequency combs, that is manifested by an array of nondegenerate, hybridized photon-magnon modes. The coupling at each node is characterized by a unique set of mode-pertaining parameters, including the wavevector Xu et al. (2024b); Christy et al. (2025), phase Joseph et al. (2024); Gardin et al. (2023), and mode profile Li et al. (2019); Hou and Liu (2019).
@@ Acknowledgments
The authors acknowledge useful discussions with Yi Li. The experimental work at UNC-CH was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences under Award Number DE-SC0026305. The theoretical work at BC was supported by U.S. National Science Foundation under Grant No. ECCS-2337713.

### v2

@@ IV Analytical Modeling
$+\sum_{m}\Big(\xi\,e^{+2\pi i\hbar f_{p}t}\,a_{+m}^{\dagger}a_{-m}^{\dagger}+\xi^{*}\,e^{-2\pi i\hbar f_{p}t}\,a_{+m}a_{-m}\Big)\,$, (5)
@@
with $\omega_{a}\equiv 2\pi f_{p}/2$ and $\Delta_{a}\equiv\pi\mathrm{FSR}$. Since each OEPO mode acts as an independent drive of magnetization dynamics, it is convenient to replace the pair-creation term in Eq. (5) with an equivalent classical drive on each mode and relabel $\omega_{\pm m}\to\omega_{\sigma m}$ for compactness, with $\sigma=\pm$ denoting the original sideband index. The Hamiltonian in Eq. (5) then becomes:
@@
$+\sum_{m,\sigma=\pm}\left(i\xi e^{-i\hbar\omega_{\sigma m}t}a_{\sigma m}^{\dagger}+\mathrm{h.c.}\right)\,.$ (7)
@@
with $g_{T}=\sqrt{g_{am}^{2}+g^{2}}$, $\kappa_{\!\mathrm{eff}}=(\kappa g_{am}^{2}+\gamma_{p}g^{2})/g_{T}^{2}$, $\kappa_{\!\mathrm{dark}}=(\kappa g^{2}+\gamma_{p}g_{am}^{2})/g_{T}^{2}$, and $\varepsilon=(\kappa-\gamma_{p})g\,g_{am}/g_{T}^{2}$. Equations (18)–(20) show that, while the dark mode $d_{\sigma m}$ couples dissipatively to $b_{\sigma m}$, it remains decoupled from the magnon mode $c_{0}$, and it is therefore invisible to the FMR probe. In the near-resonant, small-contrast limit $\lvert\kappa-\gamma_{p}\rvert\ll\kappa+\gamma_{p}$, the bright mode is well approximated as a single lossy channel with effective linewidth $\kappa_{\mathrm{eff}}$ that hybridizes with $c_{0}$ at the collective rate $g_{T}$; the corresponding two-mode cooperativity is then rewritten as $\mathcal{C}=\frac{g_{T}^{2}}{\kappa_{\!\mathrm{eff}}\,\gamma_{m}}$. Fourier transforming Eqs. (14)–(16) and solving for the MS mode response, we find
@@ V Conclusion
Third, by exploiting a dual-nonlinearity scheme, the OEMPO achieves a magnon-photon mode hybridization through combining the optoelectronic nonlinearity at the mixer (parametric pump) and the magnonic nonlinearity at the YIG (three-magnon scattering), offering a unique testbed for exploring coupled nonlinear phenomena at the hybrid interfaces among electronic, optical, and magnetic platforms Valagiannopoulos (2022). In particular, magnon modes under confined geometries or subject to nonlinear mixing processes can also give rise to a magnonic frequency comb that is controllable by external magnetic field Hula et al. (2022); Wang et al. (2024b); Christy et al. (2025). This way it is possible to achieve mode hybridization between two interfering frequency combs, that is manifested by an array of nondegenerate, hybridized photon-magnon modes. The coupling at each node is characterized by a unique set of mode-pertaining parameters, including the wavevector Xu et al. (2024b); Christy et al. (2025), phase Joseph et al. (2024); Gardin et al. (2023), and mode profile Li et al. (2019); Hou and Liu (2019).
@@ Acknowledgments
The authors acknowledge useful discussions with Yi Li. The experimental work at the UNC-CH was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences under Award Number DE-SC0026305. The theoretical work at BC was supported by U.S. National Science Foundation under Grant No. ECCS-2337713.